mathclass:

2. You complain that the hot tub in your hotel suite is not hot enough. The hotel tells you that they will increase the temperature by 10 degrees each hour. If the current temperature of the hot tub is 75ยบ F, what will be the temperature of the hot tub after 3 hours?

mathclass:

would it be 105F

jhonyy9:

using the rule of 3 simple you can write 1 hour increase by 10 degrees 3 hours increase by x degrees cross multiplie and calcule the x value so will get x = 3*10/1 = 3*10 = ? degrees

mathclass:

30

jhonyy9:

@mathclass pls dont give direct answers - this is opposite with the rules of QC

jhonyy9:

yes this is 30 degrees but the question is different

mathclass:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 @mathclass pls dont give direct answers - this is opposite with the rules of QC \(\color{#0cbb34}{\text{End of Quote}}\) I am the one who asked the question

jhonyy9:

ok so sorry

jhonyy9:

so than the temperature increase by 30 degrees in 3 hours how many will be the temperature in hotel after these 3 hours ?

mathclass:

105

jhonyy9:

yes you get it right

mathclass:

thank you

jhonyy9:

anytime just tag me i m glad to give help